5.1x - 11.4 = -0.9x - 17.4
Answer by awang1996(50)

5.1x - 11.4 = -0.9x - 17.4
+0.9x +0.9x

6x - 11.4 = -17.4
+ 11.4 +11.4
6x = -6

x = -1